How many grams of pure salt must be added to 40 grams of a 20% salt solution to make a solution that is 46% salt?
Answer by stanbon(75887)

How many grams of pure salt must be added to 40 grams of a 20% salt solution to make a solution that is 46% salt?
------------------
Equation:
salt + salt = salt
0.20(40) + 1.00x = 0.46(40+x)
Multiply thru by 100 to get:
20*40 + 100x = 46*40 + 46x
54x = 26*40
x = (26*40)/54
x = 1040/54
x = 19.26 grams of pure salt must be added
===================
Cheers,
Stan H.